using Cauchys Residue Theorem and the so-alled pacman inte 6. (10 pts) Evaluate gration contour.
Answer #1
  • \ointcf(x) dx = 2\pii { Residues of f(xn) } , where n is number of poles

If f has a removable singularity at x = x0, then the residue is equal to zero. If f has a single pole at x = x0 , then

  • Res f(x0) =  lim0 (x-x0)*f(x)

and if f has a pole of order k at x = x0, then

  • Res f(x0) = \frac{1}{(k-1)!}   lim0 dak {(x-x0)k * f(x)}  , k ∈ {1, 2, 3, . . .}

In the given problem,

\int_{0 }^{\infty }CAT 8)​​ =   \int_{0 }^{\infty }\frac{dx}{x^{1/2}(x-2(1+i))(x-2(1-i))}

Let f(x) = \frac{1}{x^{1/2}(x^{2}-4x+8)}   ,

now let us find non analytic points i.e, x2-4x+8 = 0

\Rightarrow(x-2(1+i))(x-2(1-i)) = 0

\Rightarrowx0 = 2(1+i) or x1 = 2(1-i)

The residues are,

R0 = Res f(x0) =  lim0 (x-x0)*f(x) =  lim\frac{(x-2(1+i))}{x^{1/2}(x-2(1+i))(x-2(1-i))}

=  lim  r1/2(r - 2(1- i))

  = 4i2(1 + i)

R1 = Res f(x1) =  lim1(x-x1)*f(x) =  \lim_{x\rightarrow 2(1-i)}\frac{(x-2(1-i))}{x^{1/2}(x-2(1+i))(x-2(1-i))}

=  \lim_{x\rightarrow 2(1-i)}  \frac{1}{x^{1/2}(x-2(1+i))}

  = \frac{-1}{4i \sqrt{2(1-i)}}

\int_{0 }^{\infty }f(x) dx = 2\pii { Residues of f(xn) } , where n is number of poles

\int_{0 }^{\infty }CAT 8) = 2\pii {R0 + R1}

=   2\pii * { \frac{1}{4i \sqrt{2(1+i)}}-\frac{1}{4i \sqrt{2(1-i)}} }

= 2\pii * {\frac{\sqrt{1-i}-\sqrt{1+i}}{4\sqrt{2}i * \sqrt{(1+i)(1-i)}}} =   \frac{2\pi i}{8i} * (\sqrt{1-i} - \sqrt{1+i})

(Since 1 i)(1 = \sqrt{2} )

\int_{0 }^{\infty }CAT 8) =   \frac{\pi }{4}1 i)(1